OpenStudy (anonymous):

114 percent of 95 percent is what number.

jhonyy9 (jhonyy9):

wann this being how many percent is 114 of 95 ?

jhonyy9 (jhonyy9):

or what number will be 114 percent of 95 ?

OpenStudy (anonymous):

yeaah

jhonyy9 (jhonyy9):

what from these two ?

jhonyy9 (jhonyy9):

so what wann being true ?

OpenStudy (anonymous):

Just tell me how to solve the problem in numbers though

jhonyy9 (jhonyy9):

so this depended what wann being true because not is indifferent if you wann being 114 is how many percent of 95 or what number is 95 percent of 114 ok ?

jhonyy9 (jhonyy9):

what is the 10 percent of 100 ? for example

jhonyy9 (jhonyy9):

so us the rule of 3 simple 100 is 100 percent of 100 than x will be 10 percent of 100 and now just cross multiplie and calculi the x ------------------------------------ x= 100*10 / 100 = 10 x=10 what mean that 10 percent of 100 is equal 10 ok ?

OpenStudy (anonymous):

\[ \frac{ 114 }{ 100 } * \frac{ 95 }{ 100 } = \\\ \frac{ 57}{ 50 } * \frac{ 19}{ 20 } = \\\ \frac{ 10830}{ 10000 } =1.083\]